State the chain rule, the product rule and the quotient rule of differentiation. For each rule, give an example of a trigonometric expression which can be differentiated using the rule.


1
Expert's answer
2021-05-10T13:12:28-0400

Chain rule of differentiation:


"\dfrac{dy}{dx} = \dfrac{dy}{du} \dfrac{du}{dx}"


Example: "f(x) = sin(3x^2+x)"


It looks like the outside function is the sine and the inside function is 3x2+x. The derivative is then.


"f'(x) = cos(3x^2+x)(6x+1)"


"= (6x+1)cos(3x^2+x)"


Product Rule of Differentiation:


"\dfrac{d(uv)}{dx} = u \dfrac{dv}{dx} + v \dfrac{du}{dx}"


Example : "f(x) = xsinx"


"f'(x) = x\dfrac{d(sinx)}{dx} + sinx \dfrac{d(x)}{dx}"


"f'(x) = xcosx+sinx."


Quotient Rule of differentiation:


"\dfrac{dy}{dx} = \dfrac{v\dfrac{du}{dx}-u\dfrac{dv}{dx}}{v^2}"


Example: "f(x) = \dfrac{sinx}{x}"


"f'(x) = \dfrac{x\dfrac{d(sinx)}{dx}-sinx \dfrac{d(x)}{dx}}{x^2}"


"f'(x) = \dfrac{xcosx-sinx}{x^2}"